Suppose an economy is described by following aggregate expenditure (AE) model:

C = 40 + 0.75YD    

I = 60

C is consumption (0.75 is the marginal propensity to consume)

 YD is disposable income, 

 I is investment spending.


If there is an investment boom that leads to an increase in investment of 25. Calculate the new level of GDP and explain why the old level of national income is no longer the equilibrium level.



1
Expert's answer
2021-02-16T05:50:56-0500

AE=aggregate expenditure

AE=C+I+G

At equilibrium AE=Yd

"AE=Y=40+0.75Y+60"

Expression for AE=Y at equilibrium will be

"Y=100+0.78Y"

"Y-0.75Y=100"

"0.25Y=100"

Y=400

AE"=100+0.75Y"

"=100+0.75(400)"

"=340" (the old GDP)

"Y=40+0.75Y+85"

"Y=125+0.75Y"

"Y-0.75Y=125"

"0.25Y=125"

"Y=500"

AE"=125+0.75y"

"=125+0.75(500)"

AE=415(the new GDP)

The old level of national income is no longer the equilibrium level because an increase in national income to a new one increases the level of equilibrium.